OpenStudy (anonymous):

The shoe store has twice as many black shoes as it does brown shoes. The total number of shoes is 66. How many brown shoes are there? @SomeGirl1999

OpenStudy (swissgirl):

Ok let x represent the amount of brown shoes So then there are x amount of brown shoes Now there are twice the amount of black shoes Brown shoes: x Black shoes: 2*(Brown shoes)=2*x So Black shoes + Brown shoes =66 2x+x=66 Now solve for x

OpenStudy (anonymous):

i alreday got the answer @swissgirl

OpenStudy (swissgirl):

What is it?

OpenStudy (anonymous):

22 @swissgirl

OpenStudy (amistre64):

how many purple shoes does it have?

OpenStudy (swissgirl):

OMG yaaaa black and brown shoes are borringggg Wldnt shop in that store

OpenStudy (anonymous):

i know huh what color would u buy ur shoes @swissgirl